OpenStudy (chrisplusian):

Use differentials to approximate (26)^1/3

OpenStudy (jlastino):

The given function is cube root of x Give a number that has a whole number as a cube root and is close to 26... 27 27^(1/3) = 3 So we have x1 = 27 y1= 3 (because the function is y= x^(1/3) x2= 26 y2= ? change in x = -1 Take the derivative of the function y' = (1/(2 (cube root of x^2))) x' we are looking for y' sub x=27 and x'=-1 into the equation 1/(2 (27)^(2/3)) -1 -1/2(9) y'=-1/18 Take that value then add it to 3 3- 1/18 3 - 0.06 =2.94 (answer) If you input cube root of 26 in a calcu it would give 2.96 (close enough)